# Scalable Multi-Objective Reinforcement Learning with Fairness Guarantees using Lorenz Dominance

### v1

@@ C Experiment Reproducibility Details
GPI-LS $0.30\pm 0.05$ $0.29\pm 0.11$ $0.31\pm 0.06$ $0.48\pm 0.04$-- - --- - --- - --- - --- -
PCN $0.86\pm 0.01$ $0.59\pm 0.01$ $0.65\pm 0.03$ $0.77\pm 0.01$ $0.78\pm 0.01$ $0.75\pm 0.01$ $0.48\pm 0.01$ $0.40\pm 0.01$ $0.66\pm 0.01$
LCN $\underline{0.92\pm 0.02}$ $\underline{0.64\pm 0.03}$ $0.72\pm 0.03$ $\underline{0.81\pm 0.05}$ $0.78\pm 0.02$ $0.79\pm 0.02$ $0.69\pm 0.02$ $\underline{0.78\pm 0.01}$ $0.69\pm 0.02$
LCN-Redist. $0.83\pm 0.05$ $0.59\pm 0.03$ $0.55\pm 0.03$ $0.62\pm 0.02$ $0.64\pm 0.02$ $0.59\pm 0.01$ $0.46\pm 0.01$ $0.36\pm 0.01$ $0.55\pm 0.02$
LCN-Mean $\underline{0.90\pm 0.02}$ $\underline{0.61\pm 0.02}$ $\underline{0.88\pm 0.05}$ $\underline{0.83\pm 0.04}$ $\underline{0.80\pm 0.02}$ $\underline{0.81\pm 0.02}$ $\underline{0.79\pm 0.02}$ $\underline{0.77\pm 0.01}$ $\underline{0.81\pm 0.02}$
@@
GPI-LS $0.20\pm 0.04$ $0.33\pm 0.03$ $0.40\pm 0.02$ $0.56\pm 0.02$ $0.45\pm 0.01$-- - --- - --- - --- -
PCN $0.13\pm 0.02$ $0.23\pm 0.01$ $0.32\pm 0.01$ $0.32\pm 0.01$ $0.36\pm 0.01$ $0.37\pm 0.01$ $0.45\pm 0.01$ $0.45\pm 0.00$ $0.49\pm 0.01$
LCN $0.09\pm 0.02$ $0.09\pm 0.01$ $0.12\pm 0.01$ $0.17\pm 0.02$ $0.23\pm 0.02$ $0.28\pm 0.01$ $\underline{0.23\pm 0.01}$ $\underline{0.21\pm 0.01}$ $\underline{0.19\pm 0.01}$
LCN-Redist. $\underline{0.04\pm 0.01}$ $\underline{0.04\pm 0.01}$ $\underline{0.08\pm 0.01}$ $\underline{0.15\pm 0.01}$ $\underline{0.20\pm 0.01}$ $\underline{0.20\pm 0.01}$ $0.28\pm 0.01$ $0.29\pm 0.01$ $0.35\pm 0.01$
LCN-Mean $0.07\pm 0.01$ $0.10\pm 0.02$ $0.20\pm 0.01$ $\underline{0.15\pm 0.01}$ $\underline{0.21\pm 0.02}$ $0.29\pm 0.01$ $\underline{0.21\pm 0.01}$ $0.23\pm 0.01$ $0.25\pm 0.02$
@@
GPI-LS $0.29\pm 0.09$ $0.57\pm 0.08$ $0.66\pm 0.04$ $0.67\pm 0.03$-- - --- - --- - --- - --- -
PCN $0.19\pm 0.03$ $0.45\pm 0.02$ $0.52\pm 0.02$ $0.68\pm 0.01$ $0.61\pm 0.01$ $0.59\pm 0.01$ $0.72\pm 0.01$ $0.72\pm 0.01$ $0.70\pm 0.01$
LCN $\underline{0.09\pm 0.02}$ $\underline{0.39\pm 0.02}$ $\underline{0.39\pm 0.01}$ $\underline{0.53\pm 0.04}$ $\underline{0.49\pm 0.01}$ $0.53\pm 0.01$ $\underline{0.59\pm 0.01}$ $\underline{0.69\pm 0.01}$ $\underline{0.56\pm 0.01}$
LCN-Redist. $\underline{0.10\pm 0.04}$ $0.53\pm 0.03$ $0.58\pm 0.02$ $0.61\pm 0.02$ $0.64\pm 0.01$ $0.66\pm 0.01$ $0.70\pm 0.01$ $0.71\pm 0.01$ $0.70\pm 0.01$
LCN-Mean $0.08\pm 0.02$ $0.46\pm 0.04$ $0.41\pm 0.01$ $0.56\pm 0.01$ $0.47\pm 0.01$ $\underline{0.50\pm 0.01}$ $\underline{0.61\pm 0.01}$ $\underline{0.68\pm 0.01}$ $0.59\pm 0.01$

### v2

@@ Appendix E Additional Results
GPI-LS $0.16\pm 0.03$ $0.54\pm 0.03$ $0.47\pm 0.03$ $0.32\pm 0.01$ -- -- -- -- --
PCN $0.82\pm 0.02$ $0.69\pm 0.01$ $0.42\pm 0.01$ $0.43\pm 0.01$ $0.49\pm 0.01$ $0.42\pm 0.01$ $0.34\pm 0.01$ $0.37\pm 0.01$ $0.22\pm 0.01$
LCN $0.93\pm 0.04$ $\underline{0.88\pm 0.01}$ $0.45\pm 0.07$ $\underline{0.61\pm 0.01}$ $\underline{0.66\pm 0.02}$ $0.73\pm 0.02$ $\underline{0.45\pm 0.01}$ $0.89\pm 0.01$ $0.62\pm 0.04$
LCN-Redist. $0.86\pm 0.07$ $0.71\pm 0.02$ $0.36\pm 0.03$ $0.40\pm 0.02$ $0.43\pm 0.02$ $0.41\pm 0.01$ $0.34\pm 0.01$ $0.42\pm 0.01$ $0.38\pm 0.01$
LCN-Mean $\underline{0.99\pm 0.01}$ $0.85\pm 0.02$ $\underline{0.90\pm 0.02}$ $0.54\pm 0.03$ $0.47\pm 0.02$ $\underline{0.83\pm 0.01}$ $0.39\pm 0.01$ $\underline{0.90\pm 0.01}$ $\underline{0.82\pm 0.02}$
@@
GPI-LS $\pm 0.05 0.30±0.03$$\pm 0.03$ $0.46\pm 0.02$ $0.50\pm 0.02$ $\underline{0.49\pm 0.01}$ -- -- -- --
PCN $0.10\pm 0.01$ $0.18\pm 0.01$ $0.30\pm 0.01$ $0.30\pm 0.00$ $0.30\pm 0.00$ $0.34\pm 0.00$ $0.44\pm 0.00$ $0.43\pm 0.00$ $0.46\pm 0.00$
LCN $06\pm 0.01 0.0$$09\pm 0.01$ $0.19\pm 0.01$ $0.25\pm 0.01$ $0.23\pm 0.01$ $\underline{0.25\pm 0.00}$ $\underline{0.33\pm 0.01}$ $\underline{0.30\pm 0.00}$ $\underline{0.30\pm 0.00}$
LCN-Redist. $\underline{0.01\pm 0.00}$ $\underline{0.05\pm 0.00}$ $\underline{0.18\pm 0.01}$ $\underline{0.23\pm 0.01}$ $\underline{0.21\pm 0.01}$ $0.27\pm 0.01$ $\underline{0.33\pm 0.01}$ $0.40\pm 0.01$ $0.45\pm 0.00$
LCN-Mean $0.07\pm 0.01$ $0.10\pm 0.01$ $0.24\pm 0.01$ $0.25\pm 0.01$ $0.23\pm 0.01$ $0.26\pm 0.01$ $0.39\pm 0.01$ $\underline{0.30\pm 0.00}$ $0.32\pm 0.00$
@@
GPI-LS $0.54\pm 0.05$ $0.64\pm 0.02$ $0.62\pm 0.01$ $0.70\pm 0.01$ -- -- -- -- --
PCN $0.13\pm 0.02$ $0.48\pm 0.01$ $0.58\pm 0.01$ $0.67\pm 0.01$ $3\pm 0.01 0.66±0$$6\pm 0.01$ $0.71\pm 0.01$ $0.72\pm 0.01$ $0.74\pm 0.01$
LCN $0.03\pm 0.02$ $\underline{0.37\pm 0.01}$ $\underline{0.43\pm 0.01}$ $\underline{0.52\pm 0.01}$ $\underline{0.47\pm 0.01}$ $\underline{0.50\pm 0.00}$ $\underline{0.60\pm 0.00}$ $0.67\pm 0.01$ $\underline{0.57\pm 0.01}$
LCN-Redist. $0.06\pm 0.04$ $0.45\pm 0.02$ $0.61\pm 0.02$ $0.64\pm 0.01$ $0.62\pm 0.01$ $0.64\pm 0.01$ $0.69\pm 0.01$ $0.69\pm 0.00$ $0.71\pm 0.00$
LCN-Mean $\underline{0.01\pm 0.00}$ $0.40\pm 0.02$ $0.40\pm 0.01$ $0.56\pm 0.02$ $0.54\pm 0.01$ $0.50\pm 0.01$ $0.60\pm 0.01$ $\underline{0.65\pm 0.00}$ $0.58\pm 0.00$
Gini coefficient: quantifies the reward distribution among various groups in the city. A value of 0 indicates perfect equality (equal percentage of satisfied OD flows per group), while 1 represents perfect inequality (only the demand of one group is satisfied). Although traditionally used to assess income inequality (zheng_ai_2022), it has also been employed in the context of transportation network design (feng_multicriteria_2014). In fig. 14 and table 2, we show detailed results on both Xi’an and Amsterdam for all objectives. fig. 15 and fig. 16 show the learning curves of PCN and LCN for all objectives.